OpenStudy (anonymous):

the slope of the line is -2 find the point slope. (4,-6)

OpenStudy (anonymous):

point slope from combines the slope and the point paramaters in the form of y-y(sub1)=m(x-x(sub1))

OpenStudy (anonymous):

you have m=-2, y(sub1)=-6 and x(sub1)= 4. these are your paramaters that you will plug into the equation to graph the line

OpenStudy (anonymous):

ok.....thts the thing... y-(-6)=-2(x-4) y+6=-2x+8 2x+y-2=0

OpenStudy (anonymous):

you dont set it to zero , set it to the form y=mx+b (slope intercept form)

OpenStudy (anonymous):

Here,we know tht Slope m=(y-y1)/(x-x1) here m=-2 y1=-6 x1=4

OpenStudy (anonymous):

so...the answer is y - 6 = -2(x+4) ? @ashwinjohn3

OpenStudy (anonymous):

You havent got it right... itz y-y1=m(x-x1) here y1=-6 and x1=4 m=-2 y-(-6)= -2(x-4)

OpenStudy (anonymous):

Your negatives are off from the equation :)